A hydraulic cylinder having a bore of 125 mm, a rod of 80 mm diameter and a stroke of 350mm is fully extended and retract in a total time of 15 sec. The extend thrust to be exerted by the cylinder is 20 tomes, the retract thrust being 10 tormes. Determine:


(a) The theoretical pump delivery required


(b) The theoretical system pressure when the cylinder is extending.


(c) The theoretical system pressure when the cylinder is retracting.


(d) The actual pump displacement if the volumetric efficiency is 90% and the pump drive at 1440 rev/min. (e) The maximum power input to the pump if the torque efficiency is 85%.


1
Expert's answer
2021-12-21T00:37:01-0500

system pressure when the cylinder is extending.(20 T = 196133 N)

"p = \frac{F}{\frac{\pi}{4}d^2} = \frac{196133}{\frac{\pi}{4}80^2} =156" bar

system pressure when the cylinder is retracting.(10 T = 98066 N)

"p = \frac{F}{\frac{\pi}{4}d^2} = \frac{98066}{\frac{\pi}{4}80^2} =19.50" ​bar